# Interval Selection in Sliding Windows

### v1

@@ A Lack of Lower Bounds in the Sliding Window Model
Interestingly, to the best of our knowledge, for graph problems (recall that the interval selection problem is a independent set problem on interval graphs) no separation result between the one-pass streaming and the sliding window models are known. In particular, we are not aware of any space lower bounds for graph problems specifically designed for the sliding window setting, and the only lower bounds that apply are those that carry over from the one-pass streaming setting. Our work is thus the first to establish such a separation. While our results for arbitrary-length intervals are not tight, we stress that for most problems considered, including Maximum Matching and Minimum Vertex Cover, no tight bounds are known. It is unclear whether this is due to a lack of techniques for improved algorithms or for stronger lower bounds.
@@
Our lower bound for unit-length intervals is obtained by a reduction to the $\textsf{Index}_{n}$ problem in the one-way two-party communication setting. In this setting, there are two parties, denoted Alice and Bob. Each party holds a portion of the input data. Alice sends a single message to Bob, who then outputs the result of the computation. The objective is to solve a problem using a message of smallest possible size. In $\textsf{Index}_{n}$, Alice holds a bit-string $X\in\{0,1\}^{n}$, and Bob holds an index $J\in[n]$, where $[n]=\{1,2,3,...,n\}$ and the objective for Bob is to report the bit $X[J]$. It is well-known that a message of size $\Omega(n)$ is needed to solve the problem.
@@
Despite the gap between the upper bound and the lower bound approximation factors, the result for arbitray-length interval selection is very significant. For problems related to graphs, no lower bounds in the sliding window model that are stronger than in the insertion-only model are known (e.g., for matchings, a sliding window algorithm with approximation factor 3+eps is known [7], while only the 1.69-approximation impossibility result from insertion-only streaming applies and no better impossibility results are known [13]).
@@
For Maximum Matching, a $2$-approximation with space $\tilde{\mathrm{O}}(n)$ can easily be achieved in the streaming model by running the Greedy matching algorithm, and the smooth histogram method immediately yields a $(4+\varepsilon)$-approximation sliding-window algorithm when built on Greedy. Crouch et al. [7] observed that the resulting algorithm can be analyzed more precisely and showed that it actually yields a $(3+\varepsilon)$-approximation sliding window algorithm. Regarding the weighted version of the Maximum Matching problem, the smooth histogram technique immediately yields a $(4+\varepsilon)$-approximation using the $(2+\varepsilon)$-approximation streaming algorithm by [16], and, again, as proved by Biabani et al. [3], the analysis can be tailored to the Maximum Matching problem to establish an approximation factor of $3.5+\varepsilon$ without changing the algorithm. Alexandru et al. [1] then improved the approximation factor to $3+\varepsilon$ by running the smooth histogram algorithm with a slightly different objective function.
@@ 2 Preliminaries
Every constant-error one-way communication protocol that solves $\textsf{Chain}_{k}(n)$ requires at least one message of size $\Omega(n)$.
@@ 3 Unit-length Intervals
In this section, we give our sliding window algorithm (Subsection 3.1) and our lower bound (Subsection 3.2) for unit length intervals.
@@ 3.1 Sliding Window Algorithm for Unit-length Intervals
We now describe our algorithm for unit length intervals.
@@
To see that $f$ is injective, observe that for any two intervals in $OPT$, since these intervals are independent and of unit length, the integer parts of their left endpoints are distinct. Hence, $f(I_{1})\neq f(I_{2})$, for any two distinct intervals $I_{1},I_{2}\in OPT$.
@@
To see this, observe that, for two integers $x\neq y$ of the same parity, $\texttt{latest}(x)$ and $\texttt{latest}(y)$ (if they exist) are independent. This is because $|y-x|\geq 2$ and the intervals have unit length. By the pigeonhole principle, there are at least $\frac{|\texttt{latest}|}{2}$ intervals where their indices inside latest have the same parity, which implies that $|ALG|\geq\frac{|\texttt{latest}|}{2}$.
@@ 3.2 Space Lower Bound
Let $\varepsilon>0$ be any small constant. Then, any algorithm in the sliding window model that computes a $(2-\varepsilon)$-approximate solution to Interval Selection on unit length intervals requires a memory of size $\Omega(L)$.
@@ 4.1.1 Cabello and Pérez-Lantero Algorithm
1:$\mathcal{W}\leftarrow\{\mathbb{R}\}$ a window partition
@@
2:     if there exists $W=[a,b)\in\mathcal{W}$ such that $I\subseteq W$ then
3:         if $I\cap\text{leftmost}(W)\cap\text{rightmost}(W)\neq\emptyset$ then
4:              if $\text{right}(I)\leq\text{right}(\text{leftmost}(W))$ then leftmost$(W)\leftarrow I$
5:              if $\text{left}(I)\geq\text{left}(\text{rightmost}(W))$ then rightmost$(W)\leftarrow I$
@@
7:              if $\text{right}(I)\leq\text{left}(\text{rightmost}(W))$ then
8:                  Let $W_{1}\leftarrow(a,\text{right}(I)]$
9:                  Let $W_{2}\leftarrow(\text{right}(I)),b]$
10:                  leftmost$(W_{1})\leftarrow I$, rightmost$(W_{1})\leftarrow I$
11:                  leftmost$(W_{2})\leftarrow\text{rightmost}(W)$, rightmost$(W_{2})\leftarrow\text{rightmost}(W)$
@@
13:                  Let $W_{1}\leftarrow[a,\text{left}(I))$
14:                  Let $W_{2}\leftarrow[\text{left}(I)),b)$
15:                  leftmost$(W_{2})\leftarrow I$, rightmost$(W_{2})\leftarrow I$
16:                  leftmost$(W_{1})\leftarrow\text{leftmost}(W)$,rightmost$(W_{1})\leftarrow\text{leftmost}(W)$
17:              Insert $W_{1},W_{2}$ into $\mathcal{W}$
18:              Remove $W$ from $\mathcal{W}$
@@
1:Return $\{\text{leftmost}(W)|W\in\mathcal{W}\}$
The key idea behind the algorithm is to maintain a partition $\mathcal{W}$ of the real line $\mathbb{R}$ that we refer to as a window partition. Initially, the algorithm starts with the single window $\mathcal{W}=\{\mathbb{R}\}$, and as the algorithm proceeds, the real line is partitioned into half-open intervals. This is achieved as follows. Arriving intervals that cross a window boundary are ignored. Consider thus an arriving interval $I$ that lies entirely within a window. In each window, the algorithm stores the left-most (the interval with the left-most right delimiter) and right-most (the interval with the right-most left delimiter) intervals within the window that it has observed thus far. If the interval $I$ together with either the left-most or the right-most interval of the window forms an independent set of size two then the window is split into two windows and the left-most and right-most intervals are updated accordingly. Otherwise, if $I$ intersects with both the left-most and right-most intervals of the window then $I$ is only used to potentially replace the left-most and/or right-most intervals of the window.
@@
C1 For each window $W\in\mathcal{W}$, the $\mathcal{CP}$ algorithm stores at least one (and at most two) intervals and the input instance is such that there are no two disjoint intervals that lie within window $W$. C2 The $\mathcal{CP}$ algorithm outputs a solution of size $|\mathcal{W}|$, i.e., one interval per window. Furthermore, we have that $|\mathcal{W}|\geq\frac{|OPT|+1}{2}$, i.e., the algorithm has an approximation factor slightly better than $2$. C3 The algorithm uses space $O(|OPT|)$.
@@
The output produced by the $\mathcal{CP}$ algorithm consists of one interval per window. The lemma then follows since, by construction, the number of windows cannot decrease. ∎
@@ 4.1.2 The Smooth Histogram Technique
6:       Denote by $\mathcal{CP}_{1},\mathcal{CP}_{2},\dots$ the runs sorted in increasing order with respect to their
7:          starting positions. Then, repeatedly remove a run $\mathcal{CP}_{i}$ if maintained solutions of
8:          the adjacent runs $\mathcal{CP}_{i-1}$ and $\mathcal{CP}_{i+1}$ are within a factor of $1+\varepsilon$ in size
9:     output solution of oldest run
@@
Upon the arrival of a new interval, Algorithm 3 first creates a new run of the $\mathcal{CP}$ algorithm and feeds the new interval into all currently running copies of $\mathcal{CP}$. The method relies on a clever way of deleting unnecessary runs: A run is deleted if it is expired (i.e it contains an interval which appeared before the start of the current window) and if the closest run with earlier start time and the closest run with later start time are such that their solutions differ in size by less than a $1+\varepsilon$ factor. Consider the moment after a clean-up took place, and let us denote the stored runs by $\mathcal{CP}_{1},\dots,\mathcal{CP}_{\ell}$. The clean-up rule implies that the stored runs have the properties depicted in Figure 4.
@@
Recall that the output of Algorithm 3 is the output of the oldest run of an instance of $\mathcal{CP}$, and let us denote this run by $\mathcal{CP}_{1}$. First, if the start position of $\mathcal{CP}_{1}$ coincides with the oldest interval of the current sliding window then $\mathcal{CP}_{1}$ was run on the entire window and we immediately obtain an approximation factor of $2$ (Property C2 of the $\mathcal{CP}$ algorithm). Hence, suppose that this is not the case, and denote the run that has expired most recently by $\mathcal{CP}_{0}$. Observe that, prior to $\mathcal{CP}_{0}$ expiring, the runs $\mathcal{CP}_{0}$ and $\mathcal{CP}_{1}$ were adjacent. Furthermore, the two runs differ by more than one interval since otherwise the starting position of $\mathcal{CP}_{1}$ would coincide with the oldest interval of the current sliding window. We now consider the suffix $S$ of intervals in the stream starting at the start position of run $\mathcal{CP}_{0}$. This suffix $S$ is partitioned into three parts $S=ABC$, where $A$ are the intervals that arrived prior to the starting position of $\mathcal{CP}_{1}$, $C$ are the intervals that arrived from the moment onward when the runs $\mathcal{CP}_{0}$ and $\mathcal{CP}_{1}$ became adjacent (either after a clean-up or they may have been adjacent from the moment onward when $\mathcal{CP}_{1}$ was created), and $B$ are the intervals between $A$ and $C$ (if there are any).
@@ 4.1.3 Sliding Window Algorithm
We will expand upon the smooth histogram method as described in Algorithm 4. The key idea is to exploit the structure of the windows created by the runs of $\mathcal{CP}$ in the smooth histogram algorithm. Based on these windows, we instantiate additional runs that target areas in which we expect to find many optimal intervals.
@@
Denote by $W_{1},\dots,W_{\ell}$ the windows created by $\mathcal{CP}_{i}$ thus far.
@@
    For each window $W_{i}$, we initiate a new run of $\mathcal{CP}$, i.e., this run only considers subsequent arriving intervals that lie within $W_{i}$.
@@
    For each pair of consecutive windows $W_{i},W_{i+1}$, we initiate a new run of $\mathcal{CP}$ that considers all subsequent intervals that lie within the merged window $W_{i}W_{i+1}$ (the window consisting of the left boundary of $W_{i}$ and the right boundary of $W_{i+1}$).
@@
We define $\varepsilon\geq 0$ such that
$|\mathcal{CP}(BC)|=\frac{|OPT|}{4+2\beta-\varepsilon}\ .$ (2)
In the following, we will argue that if $\varepsilon$ is close to 00 then we can find a better solution using the runs associated with $\mathcal{CP}_{1}$.
Let $\mathcal{W}$ be the windows created by $\mathcal{CP}_{0}$ at the moment when $\mathcal{CP}_{0}$ and $\mathcal{CP}_{1}$ became adjacent, i.e., the windows created by the run $\mathcal{CP}(AB)$. By Property C2, we have $\ell:=|\mathcal{W}|=|\mathcal{CP}(AB)|$. For each window $W_{i}\in\mathcal{W}$, let $x_{i}=|OPT_{C}\cap W_{i}|$, i.e., the number of optimal intervals in $C$ that lie within the window $W_{i}$. Furthermore, we define $X:=\sum_{i=1}^{\ell}x_{i}$.
In the next lemma, we prove that, provided $\varepsilon$ is small, the quantity $X$ is necessarily large, i.e., there are many optimal intervals in $C$ that lie within the windows $W_{i}$. We will later argue that the associated runs with $\mathcal{CP}_{1}$ can then be used to find many of these.
@@
$X\geq\frac{2-\varepsilon}{1+\beta}\cdot\ell\ .$
@@
Observe that $|OPT|\leq X+2\ell$, since at most $\ell$ intervals of $OPT$ can intersect the window boundaries $\mathcal{W}$, another $\ell$ intervals of $OPT_{A}\cup OPT_{B}$ can lie within the $\ell$ windows, and the remaining ones are the $X$ intervals of $OPT_{C}$.
@@
$\ell=|\mathcal{CP}(AB)|\leq(1+\beta)|\mathcal{CP}(B)|\leq(1+\beta)|\mathcal{CP}(BC)|=\frac{(1+\beta)|OPT|}{4+2\beta-\varepsilon}\leq\frac{(1+\beta)(X+2\ell)}{4+2\beta-\varepsilon}\ ,$
@@
Consider now the run $\mathcal{CP}(B)$, which coincides with the run $\mathcal{CP}_{1}$ until $\mathcal{CP}_{0}$ and $\mathcal{CP}_{1}$ became adjacent. Let $B_{1}$ be the intervals computed by this run that do not intersect the boundaries of $\mathcal{W}$ and let $B_{2}$ be the intervals computed by this run that intersect the boundaries of $\mathcal{W}$. Then, since $|B_{1}|+|B_{2}|=|\mathcal{CP}(B)|$, we have that either $|B_{1}|\geq\frac{1}{3}|\mathcal{CP}(B)|$ or $|B_{2}|\geq\frac{2}{3}|\mathcal{CP}(B)|$. We treat both cases separately in Lemmas 3 and 4:
@@
Suppose that $|B_{1}|\geq\frac{1}{3}|\mathcal{CP}(B)|$. Then, using the associated runs of $\mathcal{CP}_{1}$, we can output a solution of size at least $\frac{7-3\varepsilon}{6(1+\beta)}\ell\ .$
@@
We call a window $W_{i}$ good if it contains an interval from $B_{1}$. We output the solution obtained from the runs of $\mathcal{CP}$ on $W_{i}$, for all $i$, and if such a run on a good window leads to no intervals (i.e. $x_{i}=0$), then we output the interval from $B_{1}$ instead. Recall that $\mathcal{CP}$ outputs a solution of size $\frac{x_{i}+1}{2}$ if $x_{i}\neq 0$ (Property C2), and we stress here that the additive $+1$ is key for our analysis. We thus obtain a solution of size at least:
$S=\sum_{W_{i}\mbox{ good}}\max\left\{\frac{x_{i}+1}{2},1\right\}+\sum_{W_{i}\mbox{ bad},x_{i}\neq 0}\frac{x_{i}+1}{2}\ .$
Recall that $\sum^{\ell}_{i=1}x_{i}=X\geq\frac{2-\epsilon}{1+\beta}\cdot\ell$. Hence,
$S$ $\geq\sum_{W_{i}\mbox{ good}}\frac{x_{i}+1}{2}+\sum_{W_{i}\mbox{ bad},x_{i}\neq 0}\frac{x_{i}+1}{2}$
$\geq\frac{X+|B_{1}|}{2}$ $|B_{1}|\text{ is the number of good windows}$
$\geq\frac{2-\varepsilon}{2(1+\beta)}\ell+\frac{1}{6}|\mathcal{CP}(B)|$ By Lemma 2
$\geq\frac{2-\varepsilon}{2(1+\beta)}\ell+\frac{\ell}{6(1+\beta)}$ $|\mathcal{CP}(B)|\geq\frac{|\mathcal{CP}(AB)|}{1+\beta}=\frac{\ell}{1+\beta}\text{ by Prop. {S2}}$
$\geq\frac{7-3\varepsilon}{6(1+\beta)}\ell\ .$
@@
Suppose that $|B_{2}|\geq\frac{2}{3}|\mathcal{CP}(B)|$. Then, using the associated runs of $\mathcal{CP}_{1}$, we can output a solution of size at least $\frac{7-3\varepsilon}{6(1+\beta)}\ell\ .$
@@
Let $B_{2}^{1}\subseteq$ be the intervals of $B_{2}$ that lie on the boundary of two windows $W_{i}W_{i+1}$ where $i$ is even, and let $B_{2}^{2}=B_{2}\setminus B_{2}^{1}$. Then, either $|B_{2}^{1}|\geq\frac{1}{2}|B_{2}|$ or $|B_{2}^{2}|\geq\frac{1}{2}|B_{2}|$.

### v3

@@ A Lack of Lower Bounds in the Sliding Window Model
Interestingly, to the best of our knowledge, we are aware of only a single lower bound result for graphs in the sliding window model published before our work. Crouch et al. [7] proved an $\Omega(L)$ space lower bound for Minimum Spanning Forest in the sliding window model, while this problem can be solved with space $\tilde{\mathrm{O}}(n)$ in the streaming model.
Our work establishes such separation results for both the unweighted and the weighted versions of Interval Selection. While our results for arbitrary-length intervals are not tight, we stress that for most problems considered, including Maximum Matching and Minimum Vertex Cover, no tight bounds are known. It is unclear whether this is due to a lack of techniques for improved algorithms or for stronger lower bounds.
@@
Our lower bound for unit-length intervals is obtained by a reduction to the $\textsf{Index}_{n}$ problem in the one-way two-party communication setting. In this setting, there are two parties, denoted Alice and Bob. Each party holds a portion of the input data. Alice sends a single message to Bob, who then outputs the result of the computation. The objective is to solve a problem using a message of smallest possible size. In $\textsf{Index}_{n}$, Alice holds a bit-string $X\in\{0,1\}^{n}$, and Bob holds an index $J\in[n]$, where $[n]=\{1,2,3,...,n\}$, and the objective for Bob is to report the bit $X[J]$. It is well-known that a message of size $\Omega(n)$ is needed to solve the problem.
@@
For Maximum Matching, a $2$-approximation with space $\tilde{\mathrm{O}}(n)$ can easily be achieved in the streaming model by running the Greedy matching algorithm, and the smooth histogram method immediately yields a $(4+\varepsilon)$-approximation sliding window algorithm when built on Greedy. Crouch et al. [7] observed that the resulting algorithm can be analyzed more precisely and showed that it actually yields a $(3+\varepsilon)$-approximation sliding window algorithm. Regarding the weighted version of the Maximum Matching problem, the smooth histogram technique immediately yields a $(4+\varepsilon)$-approximation using the $(2+\varepsilon)$-approximation streaming algorithm by [15], and, again, as proved by Biabani et al. [3], the analysis can be tailored to the Maximum Matching problem to establish an approximation factor of $3.5+\varepsilon$ without changing the algorithm. Alexandru et al. [1] then improved the approximation factor to $3+\varepsilon$ by running the smooth histogram algorithm with a slightly different objective function.
@@ 2 Preliminaries
Every constant-error one-way communication protocol that solves $\textsf{Chain}_{k}(n)$ requires at least one message of size $\Omega(n/k)$.
@@ 3 Unit-length Intervals
In this section, we give our sliding window algorithm (Subsection 3.1) and our lower bound (Subsection 3.2) for unit-length intervals.
@@ 3.1 Sliding Window Algorithm for Unit-length Intervals
We now describe our algorithm for unit-length intervals.
@@
To see that $f$ is injective, observe that for any two intervals in $OPT$, since these intervals are independent and of unit-length, the integer parts of their left endpoints are distinct. Hence, $f(I_{1})\neq f(I_{2})$, for any two distinct intervals $I_{1},I_{2}\in OPT$.
@@
To see this, observe that, for two integers $x\neq y$ of the same parity, $\texttt{latest}(x)$ and $\texttt{latest}(y)$ (if they exist) are independent. This is because $|y-x|\geq 2$ and the intervals have unit-length. By the pigeonhole principle, there are at least $\frac{|\texttt{latest}|}{2}$ intervals where their indices inside latest have the same parity, which implies that $|ALG|\geq\frac{|\texttt{latest}|}{2}$.
@@ 3.2 Space Lower Bound
Let $\varepsilon>0$ be any small constant. Then, any algorithm in the sliding window model that computes a $(2-\varepsilon)$-approximate solution to Interval Selection on unit-length intervals with probability at least $2/3$ requires a memory of size $\Omega(L)$.
@@ 4.1.1 Cabello and Pérez-Lantero Algorithm
${}\mathcal{R}\leftarrow\{\mathbb{R}\}$ a region partition
@@
2:     if there exists $R=[a,b)\in\mathcal{R}$ such that $I\subseteq R$ then
3:         if $I\cap\text{leftmost}(R)\cap\text{rightmost}(R)\neq\emptyset$ then
4:              if $\text{right}(I)\leq\text{right}(\text{leftmost}(R))$ then leftmost$(R)\leftarrow I$
5:              if $\text{left}(I)\geq\text{left}(\text{rightmost}(R))$ then rightmost$(R)\leftarrow I$
@@
7:              if $\text{right}(I)\leq\text{left}(\text{rightmost}(R))$ then
8:                  Let $R_{1}\leftarrow(a,\text{right}(I)]$
9:                  Let $R_{2}\leftarrow(\text{right}(I)),b]$
10:                  leftmost$(R_{1})\leftarrow I$, rightmost$(R_{1})\leftarrow I$
11:                  leftmost$(R_{2})\leftarrow\text{rightmost}(R)$, rightmost$(R_{2})\leftarrow\text{rightmost}(R)$
@@
13:                  Let $R_{1}\leftarrow[a,\text{left}(I))$
14:                  Let $R_{2}\leftarrow[\text{left}(I)),b)$
15:                  leftmost$(R_{2})\leftarrow I$, rightmost$(R_{2})\leftarrow I$
16:                  leftmost$(R_{1})\leftarrow\text{leftmost}(R)$,rightmost$(R_{1})\leftarrow\text{leftmost}(R)$
17:              Insert $R_{1},R_{2}$ into $\mathcal{R}$
18:              Remove $R$ from $\mathcal{R}$
@@
1:Return $\{\text{leftmost}(R)|R\in\mathcal{R}\}$
The key idea behind the algorithm is to maintain a partition $\mathcal{R}$ of the real line $\mathbb{R}$ that we refer to as a region partition. Initially, the algorithm starts with the single region ${}\mathcal{R}=\{\mathbb{R}\}$, and as the algorithm proceeds, the real line is partitioned into half-open intervals. This is achieved as follows. Arriving intervals that cross a region boundary are ignored. Consider thus an arriving interval $I$ that lies entirely within a region. In each region, the algorithm stores the left-most (the interval with the left-most right delimiter) and right-most (the interval with the right-most left delimiter) intervals within the region that it has observed thus far. If the interval $I$ together with either the left-most or the right-most interval of the region forms an independent set of size two then the region is split into two regions and the left-most and right-most intervals are updated accordingly. Otherwise, if $I$ intersects with both the left-most and right-most intervals of the region then $I$ is only used to potentially replace the left-most and/or right-most intervals of the region.
@@
C1 For each reigon $R\in\mathcal{R}$, the $\mathcal{CP}$ algorithm stores at least one (and at most two) intervals and the input instance is such that there are no two disjoint intervals that lie within region $R$. C2 The $\mathcal{CP}$ algorithm outputs a solution of size $|\mathcal{R}|$, i.e., one interval per region. Furthermore, we have that $|\mathcal{R}|\geq\frac{|OPT|+1}{2}$, i.e., the algorithm has an approximation factor slightly better than $2$. C3 The algorithm uses space $O(|OPT|)$.
@@
The output produced by the $\mathcal{CP}$ algorithm consists of one interval per region. The lemma then follows since, by construction, the number of regions cannot decrease. ∎
@@ 4.1.2 The Smooth Histogram Technique
6:       Let $\mathcal{CP}_{1},\mathcal{CP}_{2},\dots$ be the runs sorted in increasing order with respect to their starting positions.
7:       Fix $i=1$.
8:       While $\mathcal{CP}_{i}$ exists, do the following:
9:          Find maximal $j>i$ such that $|\mathcal{CP}_{i}|\leq(1+\varepsilon)|\mathcal{CP}_{j}|$
10:          If such $j$ exists, delete all the instances of $\mathcal{CP}$ strictly between $i$ and $j$.
11:          Renumber the runs sorted in increasing order with respect to their starting positions.
12:          Increase $i$ by 1.
13:     output solution of oldest run
@@
Upon the arrival of a new interval, Algorithm 3 first creates a new run of the $\mathcal{CP}$ algorithm and feeds the new interval into all currently running copies of $\mathcal{CP}$. The method relies on a clever way of deleting unnecessary runs. First, we delete a potentially expired run, i.e., a run which contains an interval that appeared before the start of the current sliding window. Then, for each run $\mathcal{CP}_{i}$, we take the latest run $\mathcal{CP}_{j}$ such that $|\mathcal{CP}_{i}|\leq(1+\varepsilon)|\mathcal{CP}_{j}|$, provided such $j>i$ exists, and we delete every run appearing later than $\mathcal{CP}_{i}$ but earlier than $\mathcal{CP}_{j}$.
Consider the moment after a clean-up took place, and let us denote the stored runs by $\mathcal{CP}_{1},\dots,\mathcal{CP}_{\ell}$. The clean-up rule implies that the stored runs have the properties depicted in Figure 4.
@@
Recall that the output of Algorithm 3 is the output of the oldest run of an instance of $\mathcal{CP}$, and let us denote this run by $\mathcal{CP}_{1}$. First, if the start position of $\mathcal{CP}_{1}$ coincides with the oldest interval of the current sliding window then $\mathcal{CP}_{1}$ was run on the entire region and we immediately obtain an approximation factor of $2$ (Property C2 of the $\mathcal{CP}$ algorithm). Hence, suppose that this is not the case, and denote the run that has expired most recently by $\mathcal{CP}_{0}$. Observe that, prior to $\mathcal{CP}_{0}$ expiring, the runs $\mathcal{CP}_{0}$ and $\mathcal{CP}_{1}$ were adjacent. Furthermore, the two runs differ by more than one interval since otherwise the starting position of $\mathcal{CP}_{1}$ would coincide with the oldest interval of the current sliding window. We now consider the suffix $S$ of intervals in the stream starting at the start position of run $\mathcal{CP}_{0}$. This suffix $S$ is partitioned into three parts $S=ABC$, where $A$ are the intervals that arrived prior to the starting position of $\mathcal{CP}_{1}$, $C$ are the intervals that arrived from the moment onward when the runs $\mathcal{CP}_{0}$ and $\mathcal{CP}_{1}$ became adjacent (either after a clean-up or they may have been adjacent from the moment onward when $\mathcal{CP}_{1}$ was created), and $B$ are the intervals between $A$ and $C$ (if there are any).
@@ 4.1.3 Sliding Window Algorithm
We will expand upon the smooth histogram method as described in Algorithm 4. The key idea is to exploit the structure of the regions created by the runs of $\mathcal{CP}$ in the smooth histogram algorithm. Based on these regions, we instantiate additional runs that target areas in which we expect to find many optimal intervals.
@@
Denote by $R_{1},\dots,R_{\ell}$ the regions created by $\mathcal{CP}_{i}$ thus far.
@@
    For each region $R_{i}$, we initiate a new run of $\mathcal{CP}$, i.e., this run only considers subsequent arriving intervals that lie within $R_{i}$.
@@
    For each pair of consecutive regions $R_{i},R_{i+1}$, we initiate a new run of $\mathcal{CP}$ that considers all subsequent intervals that lie within the merged region $R_{i}R_{i+1}$ (the region consisting of the left boundary of $R_{i}$ and the right boundary of $R_{i+1}$).
@@
We define $\varepsilon^{\prime}\geq 0$ such that
$|\mathcal{CP}(BC)|=\frac{|OPT|}{4+2\beta-\varepsilon^{\prime}}\ .$ (2)
In the following, we will argue that if $\varepsilon^{\prime}$ is close to 00 then we can find a better solution using the runs associated with $\mathcal{CP}_{1}$.
Let $\mathcal{R}$ be the regions created by $\mathcal{CP}_{0}$ at the moment when $\mathcal{CP}_{0}$ and $\mathcal{CP}_{1}$ became adjacent, i.e., the regions created by the run $\mathcal{CP}(AB)$. By Property C2, we have $\ell:=|\mathcal{R}|=|\mathcal{CP}(AB)|$. For each region $R_{i}\in\mathcal{R}$, let $x_{i}=|OPT_{C}\cap R_{i}|$, i.e., the number of optimal intervals in $C$ that lie within the region $R_{i}$. Furthermore, we define $X:=\sum_{i=1}^{\ell}x_{i}$.
In the next lemma, we prove that, provided $\varepsilon^{\prime}$ is small, the quantity $X$ is necessarily large, i.e., there are many optimal intervals in $C$ that lie within the regions $R_{i}$. We will later argue that the associated runs with $\mathcal{CP}_{1}$ can then be used to find many of these.
@@
$X\geq\frac{2-\varepsilon^{\prime}}{1+\beta}\cdot\ell\ .$
@@
Observe that $|OPT|\leq X+2\ell$, since at most $\ell$ intervals of $OPT$ can intersect the region boundaries $\mathcal{R}$, another $\ell$ intervals of $OPT_{A}\cup OPT_{B}$ can lie within the $\ell$ regions, and the remaining ones are the $X$ intervals of $OPT_{C}$.
@@
$\ell=|\mathcal{CP}(AB)|\leq(1+\beta)|\mathcal{CP}(B)|\leq(1+\beta)|\mathcal{CP}(BC)|=\frac{(1+\beta)|OPT|}{4+2\beta-\varepsilon^{\prime}}\leq\frac{(1+\beta)(X+2\ell)}{4+2\beta-\varepsilon^{\prime}}\ ,$
@@
Consider now the run $\mathcal{CP}(B)$, which coincides with the run $\mathcal{CP}_{1}$ until $\mathcal{CP}_{0}$ and $\mathcal{CP}_{1}$ became adjacent. Let $B_{1}$ be the intervals computed by this run that do not intersect the boundaries of $\mathcal{R}$ and let $B_{2}$ be the intervals computed by this run that intersect the boundaries of $\mathcal{R}$. Then, since $|B_{1}|+|B_{2}|=|\mathcal{CP}(B)|$, we have that either $|B_{1}|\geq\frac{1}{3}|\mathcal{CP}(B)|$ or $|B_{2}|\geq\frac{2}{3}|\mathcal{CP}(B)|$. We treat both cases separately in Lemmas 3 and 4:
@@
Suppose that $|B_{1}|\geq\frac{1}{3}|\mathcal{CP}(B)|$. Then, using the associated runs of $\mathcal{CP}_{1}$, we can output a solution of size at least $\frac{7-3\varepsilon^{\prime}}{6(1+\beta)}\ell\ .$
@@
We call a region $R_{i}$ good if it contains an interval from $B_{1}$. We output the solution obtained from the runs of $\mathcal{CP}$ on $R_{i}$, for all $i$, and if such a run on a good region leads to no intervals (i.e. $x_{i}=0$), then we output the interval from $B_{1}$ instead. Recall that $\mathcal{CP}$ outputs a solution of size $\frac{x_{i}+1}{2}$ if $x_{i}\neq 0$ (Property C2), and we stress here that the additive $+1$ is key for our analysis. We thus obtain a solution of size at least:
$S=\sum_{R_{i}\mbox{ good}}\max\left\{\frac{x_{i}+1}{2},1\right\}+\sum_{R_{i}\mbox{ bad},x_{i}\neq 0}\frac{x_{i}+1}{2}\ .$
Recall that $\sum^{\ell}_{i=1}x_{i}=X\geq\frac{2-\varepsilon^{\prime}}{1+\beta}\cdot\ell$. Hence,
$S$ $\geq\sum_{R_{i}\mbox{ good}}\frac{x_{i}+1}{2}+\sum_{R_{i}\mbox{ bad},x_{i}\neq 0}\frac{x_{i}+1}{2}$
$\geq\frac{X+|B_{1}|}{2}$ $|B_{1}|\text{ is the number of good regions}$
$\geq\frac{2-\varepsilon^{\prime}}{2(1+\beta)}\ell+\frac{1}{6}|\mathcal{CP}(B)|$ By Lemma 2
$\geq\frac{2-\varepsilon^{\prime}}{2(1+\beta)}\ell+\frac{\ell}{6(1+\beta)}$ $|\mathcal{CP}(B)|\geq\frac{|\mathcal{CP}(AB)|}{1+\beta}=\frac{\ell}{1+\beta}\text{ by Prop. {S2}}$
$\geq\frac{7-3\varepsilon^{\prime}}{6(1+\beta)}\ell\ .$
@@
Suppose that $|B_{2}|\geq\frac{2}{3}|\mathcal{CP}(B)|$. Then, using the associated runs of $\mathcal{CP}_{1}$, we can output a solution of size at least $\frac{7-3\varepsilon^{\prime}}{6(1+\beta)}\ell\ .$
@@
Let $B_{2}^{1}\subseteq$ be the intervals of $B_{2}$ that lie on the boundary of two regions $R_{i}R_{i+1}$ where $i$ is even, and let $B_{2}^{2}=B_{2}\setminus B_{2}^{1}$. Then, either $|B_{2}^{1}|\geq\frac{1}{2}|B_{2}|$ or $|B_{2}^{2}|\geq\frac{1}{2}|B_{2}|$.